OpenStudy (anonymous):

How do you simplify a radical on top of a radical if both radicals are not perfect squares??

OpenStudy (anonymous):

You can multiply the bottom and top by whatever is on the bottom, which will rationalize your denominator.

OpenStudy (btaylor):

or do this: \[\sqrt{x}/\sqrt{y}\rightarrow \sqrt{x/y}\] It might be able to help you.

OpenStudy (anonymous):

That also =)

OpenStudy (anonymous):

so u look at them as separate radicals?

OpenStudy (anonymous):

You can split them or join them as you like, since they have the same exponent.